OpenStudy (anonymous):

Which of the following is not necessary for a pair of angles to be vertical angles? A. They are nonadjacent B. They are congruent C. They have the same vertex D. They are complementary

OpenStudy (campbell_st):

do you know what vertical angles are...?

OpenStudy (campbell_st):

here is a website that will explain them http://www.mathopenref.com/anglesvertical.html

OpenStudy (anonymous):

i think it is A not sure

OpenStudy (campbell_st):

adjacent means the angles are next to each other... vertical angles means that they are on opposite sides of a vertex... so they must be non adjacent.... (they can't be next to each other) so it only leave 1 choice

OpenStudy (anonymous):

 A. 1 B. 2 C. 3 D. Cannot be determined

OpenStudy (campbell_st):

the word complementary means the angles add to 90 degrees... the only time they can be complementary and vertical is if they measure 45 degrees.

OpenStudy (campbell_st):

if the angles in the next question are equal x + 1 = 4x - 2 you need to solve the equation for x...

OpenStudy (anonymous):

would it b 3

OpenStudy (radar):

Mystery to solve. If x + 1 = 4x -2..................would it be 3? Solve mystery to see if it is 3. 3 + 1 = 4(3) - 2 4 =12 -2 4 = 10 Noooo it ain't 3.

OpenStudy (radar):

Now too work it using algebra. x+1=4x-2 subtract x from both sides 1 = 3x -2 add 2 to both sides. 3 =3x divide both sides by 3 1 = x x =1 check 1 + 1= 4 -2 2 = 2 checks out. mystery solved.